# Multilevel lattice codes from Hurwitz quaternion integers

### v2

@@ V-C Poltyrev-goodness
We will see now that Construction $\pi_{A}$ lattices belong to the class of lattices obtained by reduction. The case of Construction $\pi_{A}$ lattices for quaternions will be the focus of our analysis, but similar conclusions can be performed for Construction $\pi_{A}$ lattices over $\mathbb{Z},\mathbb{Z}[i]$ and $\mathbb{Z}[\omega]$. This allows to extend our next result, Proposition 3, also to these cases already approached in [1] using a different technique.
The division ring of quaternions $\mathbb{H}$ can be identified with $\mathbb{R}^{4}$ using the natural mapping
@@
By considering this mapping, the Hurwitz integers corres-pond to $\sigma^{-1}(\mathscr{H})=D_{4}^{*}$, the checkerboard $\mathbb{Z}$-lattice $D_{4}$ dual in dimension 4 [16]. Moreover, for any rational prime $p$, there exists an isomorphism $\mathscr{H}/p\mathscr{H}\sim\text{M}_{2}(\mathbb{F}_{p})$, where $\text{M}_{2}(\mathbb{F}_{p})$ represents the ring of matrices with entries in $\mathbb{F}_{p}$.
Such a ring isomorphism can be achieved by
@@
$x_{0}+x_{1}i+x_{2}j+x_{3}k$ $\mapsto\begin{pmatrix}x_{0}+x_{2}a-x_{3}b&-x_{1}+x_{2}b+x_{3}a\\ x_{1}+x_{2}b+x_{3}a&x_{0}-x_{2}a+x_{3}b\end{pmatrix},$ (18)
where $a^{2}+b^{2}+1\equiv 0\pmod{p}$, [51].
In this way, composing these two mappings enables us to obtain a reduction from a quaternionic lattice, specifically from Hurwitz lattices, to the ring of matrices $\text{M}_{2}(\mathbb{F}_{p})$.
$\phi_{p}^{\mathscr{H}}:D_{4}^{*}$ $\stackrel{{\scriptstyle\sigma}}{{\longrightarrow}}\mathscr{H}/p\mathscr{H}\stackrel{{\scriptstyle\rho}}{{\longrightarrow}}\text{M}_{2}(\mathbb{F}_{p})$
Therefore,
$\Lambda_{p}^{\mathscr{H}}(\mathscr{C})=(\phi_{p}^{\mathscr{H}})^{-1}(\mathscr{C})$
@@
Consider the case where $p=\text{Nrm}(\pi)$, $p$ is an odd rational prime. We can choose codes $\mathscr{C}^{(1)}$ and $\mathscr{C}^{(2)}$ from $\mathscr{H}/\mathscr{H}\pi$ and $\mathscr{H}/\mathscr{H}\overline{\pi}$, respectively. By applying Construction $\pi_{A}$ over Hurwitz integers (Definition 12), we obtain a code $\mathscr{C}$ in $\mathscr{H}/p\mathscr{H}\cong\text{M}_{2}(\mathbb{F}_{p})$. Thus, we have
$\phi_{\pi_{A}}^{\mathscr{H}_{p}}:D_{4}^{*}$ $\stackrel{{\scriptstyle\sigma}}{{\longrightarrow}}\mathscr{H}/p\mathscr{H}\stackrel{{\scriptstyle\varphi}}{{\longrightarrow}}\mathscr{H}/\mathscr{H}\pi\times\mathscr{H}/\mathscr{H}\overline{\pi}$
Suppose $\mathscr{C}^{(1)}$ and $\mathscr{C}^{(2)}$ are linear codes over $\mathscr{H}/\mathscr{H}\pi$ and $\mathscr{H}/\mathscr{H}\overline{\pi}$, respectively. By abuse of notation, we will also denote by $\phi_{\pi_{A}}^{\mathscr{H}_{p}}$ the reduction applied componentwise to a vector in $\mathscr{H}^{n}\cong(D_{4}^{*})^{n}$. In that case,
$(\phi_{\pi_{A}}^{\mathscr{H}_{p}})^{-1}(\mathscr{C}^{(1)},\mathscr{C}^{(2)})=\Lambda_{\pi_{A}}^{\mathscr{H}_{p}}(\mathscr{C})\subset\mathbb{R}^{4n}$
represents a Hurwitz Construction $\pi_{A}$ lattice.
Moreover, considering the reduction $\phi_{\pi_{A}}^{\mathscr{H}_{p}}:(D_{4}^{*})^{n}\rightarrow(\mathscr{H}/p\mathscr{H})^{n}$. Since $p\in\mathbb{Z}$ is an odd rational prime number, we have,
$\text{vol}((p\Lambda)^{n})=$ $\text{Nrm}(p)^{2n}\text{vol}(\Lambda)^{n}=p^{4n}(1/2)^{n}$
$d_{\text{min}}((p\Lambda)^{n})=$ $\sqrt{\text{Nrm}(p)}d_{\text{min}}(\Lambda)=p$
Therefore, to verify the non-degeneracy condition (Remark 3 in [50]), we have $\ker({\phi_{\pi_{A}}^{\mathscr{H}_{p}}})=(p\Lambda)^{n}\subseteq\mathbb{R}^{m}$ then $4n\leq m$, and,
$\liminf_{p\rightarrow\infty}{\dfrac{d_{\text{min}}((p\Lambda)^{n})}{\text{vol}((p\Lambda)^{n})}}$ $=\liminf_{p\rightarrow\infty}{\dfrac{p}{[p^{4}(1/2)]^{n/m}}}$
$=\liminf_{p\rightarrow\infty}{p^{\frac{m-4n}{m}}2^{n/m}}>0$
where $m\geq 4n\Rightarrow\frac{m-4n}{m}\geq 0$.
Let $\mathscr{C}_{k}$ be a balanced set of Construction $\pi_{A}$ codes, and let $\mathbb{L}_{p}$ be the associated lattice ensembles defined as follows
$\mathbb{L}_{p}=\left\{\Lambda_{p}=\beta\Lambda_{\pi_{A}}^{\mathscr{H}_{p}}(\mathscr{C}):\mathscr{C}\in\mathscr{C}_{k}\right\}.$ (19)
With these definitions we have the following adapted version from Minkowski-Hlawka Theorem, [14], [50], [52], for Construction $\pi_{A}$ over Hurwitz integers case.
@@
Let $(p_{j})_{j=1}^{\infty}$ be an increasing sequence of odd primes and $\mathscr{C}_{k}$ be a balanced set of codes with rank $k>n/2$. Consider the ensemble of lattices in $\mathbb{R}^{4n}$,
$\mathbb{L}_{p_{j}}=\{\Lambda_{p_{j}}=\beta\Lambda_{\pi_{A}}^{\mathscr{H}_{p_{j}}}(\mathscr{C}):\mathscr{C}\in\mathscr{C}_{k}\},$
where $\beta$ represents a normalization factor ensuring that all lattices in $\mathbb{L}_{p_{j}}$ have the same volume $V$. Let $f$ be a semi-admissible function [^2] A Riemann-integrable function $f$ is considered semi-admissible if $|f(\mathbf{x})|\leq\dfrac{b}{(1+|\mathbf{x}|)^{m+\delta}},\forall\mathbf{x}\in\mathscr{D}_{f}$, where $b>0$ and $\delta>0$ are positive constants.. If we have the non-degeneracy condition,
[…]
Then,
@@
VI Conclusion and Perspectives

### v3

@@ V-C Poltyrev-goodness
Our objective is to show Construction $\pi_{A}$ lattices over Hurwitz integers that satisfy the requirements of the Minkowski-Hlawka Theorem. This consists of generating an ensemble of lattices using codes from a balanced set. Additionally, the reduction map used to obtain the lattice must meet a non-degeneracy condition.
We will see first, a class of Construction $\pi_{A}$ lattices which belong to the set of lattices obtained by reduction [53]. The focus of our analysis is the isomorphism of Theorem 2. Similar conclusions can be derived for Construction $\pi_{A}$ lattices over $\mathbb{Z},\mathbb{Z}[i]$ and $\mathbb{Z}[\omega]$. This allows to extend the next result, Proposition 3, also to these cases already approached in [3] using a different technique.
The division ring of quaternions $\mathbb{H}$ can be identified with $\mathbb{R}^{4}$ using the natural map
@@
By considering this map, the Hurwitz integers corresponds to $D_{4}^{*}$, $\sigma^{-1}(\mathscr{H})=D_{4}^{*}$, the dual of the checkerboard $\mathbb{Z}$-lattice $D_{4}$ in dimension 4, [15]. Moreover, for any rational odd prime $p$, there exists an isomorphism $\mathscr{H}/p\mathscr{H}\cong\text{M}_{2}(\mathbb{F}_{p})$, where $\text{M}_{2}(\mathbb{F}_{p})$ represents the ring of $2\times 2$ matrices with entries in $\mathbb{F}_{p}$.
Such a ring isomorphism can be defined as
@@
$x_{0}+x_{1}i+x_{2}j+x_{3}k$ $\mapsto\begin{pmatrix}x_{0}+x_{2}a-x_{3}b&-x_{1}+x_{2}b+x_{3}a\\ x_{1}+x_{2}b+x_{3}a&x_{0}-x_{2}a+x_{3}b\end{pmatrix},$
where $a,b$ are integers and $a^{2}+b^{2}+1\equiv 0\pmod{p}$, [54].
In this way, composing these mappings enables us to obtain a reduction (surjective homomorphism) from $D_{4}^{*}$ to the ring of matrices $\text{M}_{2}(\mathbb{F}_{p})$. By considering,
$\phi_{p}^{\mathscr{H}}:D_{4}^{*}$ $\stackrel{{\scriptstyle\sigma}}{{\longrightarrow}}\mathscr{H}\stackrel{{\scriptstyle\text{mod}p}}{{\longrightarrow}}\mathscr{H}/p\mathscr{H}\stackrel{{\scriptstyle\rho}}{{\longrightarrow}}\text{M}_{2}(\mathbb{F}_{p}),$
and, given a linear code $\mathscr{C}\subset M_{2}(\mathbb{K}_{p})$,
$\Lambda_{p}^{\mathscr{H}}(\mathscr{C})=(\phi_{p}^{\mathscr{H}})^{-1}(\mathscr{C}),$
@@
Consider the case where $p=\text{Nrm}(\pi)$ is an odd rational prime. We can choose $\mathscr{C}^{(1)}$ and $\mathscr{C}^{(2)}$ linear codes from $(\mathscr{H}/\mathscr{H}\pi)^{n}$ and $(\mathscr{H}/\mathscr{H}\overline{\pi})^{n}$, respectively, then $\mathscr{C}\subset(\mathscr{H}/p\mathscr{H})^{n}$. By abuse of notation, we will also denote by $\phi_{p}^{\mathscr{H}}$ the reduction applied component-wise to a vector in $\mathscr{H}^{n}\cong(D_{4}^{*})^{n}$. Therefore, for Construction $\pi_{A}$ codes over Hurwitz integers (Definition 12), we have,
$(\phi_{p}^{\mathscr{H}})^{-1}(\mathscr{C})=\Lambda_{\pi_{A}}^{\mathscr{H}_{p}}(\mathscr{C})\subset\mathbb{R}^{4n}.$
Let $\mathscr{C}_{b_{k}}$ be the balanced set of Construction $\pi_{A}$ codes from Lemma 1, i.e, considering $\mathscr{C}^{(1)}\in\mathscr{C}_{b_{k_{1}}}$ and $\mathscr{C}^{(2)}\in\mathscr{C}_{b_{k_{2}}}$ with $\mathscr{C}_{b_{k_{1}}},\mathscr{C}_{b_{k_{2}}}$ the balanced set of all codes of rank $k_{1},k_{2}$, respectively and $k=\max\{k_{1},k_{2}\}$. Then, let $\mathbb{L}_{p}$ be the associated lattice ensembles defined as
$\mathbb{L}_{p}=\left\{\Lambda_{p}=\beta\Lambda_{\pi_{A}}^{\mathscr{H}_{p}}(\mathscr{C}):\mathscr{C}\in\mathscr{C}_{b_{k}}\right\}.$ (16)
We also remark that, since $\ker(\phi_{p_{j}}^{\mathscr{H}})=p\mathscr{H}$ [^3] Here, $\ker$ denotes the kernel of the linear map $\phi_{p}^{\mathscr{H}}$, that is, $\ker(\phi_{p}^{\mathscr{H}})=\{\lambda\in D_{4}^{*};\phi_{p}^{\mathscr{H}}(\lambda)=0\}$., it follows that
$d_{\text{min}}(\ker(\phi_{p_{j}}^{\mathscr{H}}))=p>p^{1/2},$
and the non-degeneracy condition requested in [53] is valid for any prime number. Therefore, the following version from Minkowski-Hlawka Theorem, [27, 53, 55], holds for the Construction $\pi_{A}$ over Hurwitz integers described above.
@@
Let $(p_{j})_{j=1}^{\infty}$ be an increasing sequence of odd primes and $\mathscr{C}_{b_{k}}$ be a balanced set of Construction $\pi_{A}$ codes with rank $k>n/2$. Consider the ensemble of lattices in $\mathbb{R}^{4n}$,
$\mathbb{L}_{p_{j}}=\{\Lambda_{p_{j}}=\beta\Lambda_{\pi_{A}}^{\mathscr{H}_{p_{j}}}(\mathscr{C}):\mathscr{C}\in\mathscr{C}_{b_{k}}\},$
where $\beta$ represents a normalization factor ensuring that all lattices in $\mathbb{L}_{p_{j}}$ have the same volume $V$. Let $f$ be a semi-admissible function [^4] A Riemann-integrable function $f$ is considered semi-admissible if $|f(x)|\leq\dfrac{b}{(1+|x|)^{m+\delta}},\forall x\in\mathscr{D}_{f}$, where $b>0$ and $\delta>0$ are positive constants.. If we have the non-degeneracy condition,
$d_{\text{min}}(\ker(\phi_{p_{j}}^{\mathscr{H}}))\geq c\cdot p_{j}^{1/2},$
for some constant $c>0$. Then,
@@
VI Application to Lattice Index Coding
Lattice index coding [21] is a communication problem that exploits the presence of side information at receivers to reduce decoding complexity and improve performance over an AWGN broadcast channel. It involves encoding multiple messages into lattice codewords in such way that the receivers with some knowledge (side information) can decode more efficiently.
Consider $r$ independent messages $w_{1},\ldots,w_{r}$ each from a finite set $W_{1},\ldots,W_{r}$ jointly encoded into a codeword $x=f(w_{1},\ldots,w_{r})\in\mathscr{C}$, where $\mathscr{C}\subset\mathbb{R}^{n}$ is an $n$-dimensional constellation derived from a lattice. The received signal at receiver $l\in\{1,\ldots,L\}$ is
$y_{l}=x+z_{l},\quad z_{l}\sim N(0,\sigma^{2}).$
Each receiver $l$ is characterised by a pair $(\mathrm{SNR},S_{l})$, where $\mathrm{SNR}$ is the signal-to-noise ratio and $S_{l}\subset\{1,\dots,r\}$ denotes the subset of messages known as side information and requires all messages $\{w_{1},\ldots,w_{r}\}$. Each message $w_{j}$ is transmitted at a rate $R_{j}=\frac{1}{n}\log_{2}|W_{j}|$ (in bits/dim), and the total rate is $R=\sum_{j=1}^{r}R_{j}$. The presence of side information $w_{S_{l}}=\{w_{j}:j\in S_{l}\}$ at a receiver reduces the SNR required to achieve the capacity region for the AWGN network [56], since a receiver without side information requires an SNR of $\approx 2^{2R}$, while one with side information of rate $R_{S_{l}}=\sum_{j\in S_{l}}R_{j}$ only requires $\approx 2^{2(R-R_{S_{l}})}$, [21, 42].
This improvement is observed as an increase in the minimum squared Euclidean distance between valid codewords. Let $d_{0}$ denote the minimum distance of the code without side information, and $d_{S_{l}}$ that with side information set $S_{l}$, then the squared distance gain is given by $10\log_{10}(d_{S_{l}}^{2}/d_{0}^{2})$ dB. The side information gain of a code $\mathscr{C}$ for a index set $S_{l}$ is defined as [21]
[…]
and measure the SNR gain per bit/dim of side information, relative to the baseline code performance.
Construction $\pi_{A}$ lattices is well suited for designing lattice index codes, particularly when the code is built from structured components over algebraic integer rings such as Hurwitz integers.
As an example, let us construct a code over the $\mathscr{H}$ using the primes $p_{1}=3$ and $p_{2}=5$, both of which splits in $\mathscr{H}$, $3=(1+i+j)(1-i-j)=\pi_{1}\cdot\overline{\pi}_{1}\quad\text{and}\quad 5=(1+2i)(1-2i)=\pi_{2}\cdot\overline{\pi}_{2},$ and let $q=3\cdot 5=15$. Construction $\pi_{A}$ over $\mathscr{H}$ then yields the ring isomorphism,
$\varphi^{-1}:\mathscr{H}/\mathscr{H}\pi_{1}\times\mathscr{H}/\mathscr{H}\overline{\pi}_{1}\times\mathscr{H}/\mathscr{H}\pi_{2}\times\mathscr{H}/\mathscr{H}\overline{\pi}_{2}\cong\mathscr{H}/15\mathscr{H}.$
This allows to assemble a code $\mathscr{C}\subset\mathscr{H}/15\mathscr{H}$ from component codes $\mathscr{C}_{i}\subset\mathscr{H}/\mathscr{H}\pi_{i}$ or $\mathscr{H}/\mathscr{H}\overline{\pi}_{i}$. In our example, we relabel the component codes as $\mathscr{C}_{1},\mathscr{C}_{2},\mathscr{C}_{3}$ and $\mathscr{C}_{4}$ and encode messages $a_{i}\in\mathscr{C}_{i},i=1,\ldots,4$ into the codeword,
$x$ $=\varphi^{-1}(a_{1},a_{2},a_{3},a_{4})$
$=(5\overline{\pi}_{1}a_{1}+5\pi_{1}a_{2}+3\overline{\pi}_{2}a_{3}+3\pi_{2}a_{4})\mod 15\mathscr{H}.$
Let, $\mathscr{C}_{1}=\langle\frac{1-i+j+k}{2},\frac{-1-i+j-k}{2}\rangle,\mathscr{C}_{2}=\langle 1+k\rangle,\mathscr{C}_{3}=\langle\frac{1+i+j+k}{2}\rangle$ and $\mathscr{C}_{4}=\langle\frac{-1+3i-j+3k}{2}\rangle$, linear codes over $\mathscr{H}/\mathscr{H}\pi_{1},$ $\mathscr{H}/\mathscr{H}\overline{\pi}_{1},$ $\mathscr{H}/\mathscr{H}\pi_{2}$ and $\mathscr{H}/\mathscr{H}\overline{\pi}_{2}$, respectively. Since $|\mathscr{C}_{1}|=9$, $|\mathscr{C}_{2}|=3$ and $|\mathscr{C}_{3}|=|\mathscr{C}_{4}|=5$, the total code size is $|\mathscr{C}|=3^{3}\cdot 5^{2}=675$, and as each codeword lies in a 4-dimensional real space, the rate of each message is,
$R_{1}=\frac{1}{2}\log_{2}3,\quad R_{2}=\frac{1}{4}\log_{2}3,$
$R_{3}=\frac{1}{4}\log_{2}5\quad\text{and}\quad R_{4}=\frac{1}{4}\log_{2}5.$
With no side information, the receiver decodes to the nearest point in $\mathscr{C}$, which has minimum Euclidean distance $d_{0}=1$. Now, suppose the receiver knows the first information $a_{1}=0$, corresponding to side information $S=\{1\}$. The resulting subcode is,
$\mathscr{C}_{S}=\{5\overline{\pi}_{1}\cdot 0+5\pi_{1}a_{2}+3\overline{\pi}_{2}a_{3}+3\pi_{2}a_{4}\mod 15\mathscr{H};$
$a_{j}\in\mathscr{C}_{j},j\in S^{c}\},$
which has size $|\mathscr{C}_{S}|=75$, side information rate of $R_{S}=\frac{1}{2}\log_{2}3$ and minimum distance $d_{S}=3$. The side information gain is given by,
$\Gamma(\mathscr{C},S)=\dfrac{10\log_{10}(\frac{3}{1})^{2}}{\frac{1}{2}\log_{2}3}\approx 12.0412.$
Now considering $S=\{1,2\}$, meaning the receiver knows $a_{1}=a_{2}=0$. The subcode becomes,
$\mathscr{C}_{S}=\{5\overline{\pi}_{1}\cdot 0+5\pi_{1}\cdot 0+3\overline{\pi}_{2}a_{3}+3\pi_{2}a_{4}\mod 15\mathscr{H};$
$a_{j}\in\mathscr{C}_{j},j\in S^{c}\},$
with minimum distance $d_{S}=3$, $|\mathscr{C}_{S}|=25$ and the side information rate $R_{S}=R_{1}+R_{2}=\frac{1}{4}\log_{2}9+\frac{1}{4}\log_{2}3=\frac{1}{4}\log_{2}27$.
In general, for any side information set $S\subset\{1,2,3,4\}$ the side information rate is given by $R_{S}=\frac{1}{4}\log_{2}\prod_{j\in S}|\mathscr{C}_{j}|$.
The side information gain in this case is,
$\Gamma(\mathscr{C},S)=\dfrac{10\log_{10}(\frac{3}{1})^{2}}{\frac{1}{4}\log_{2}27}\approx 8.0274.$
By repeating this same analysis for all subsets $S\subset\{1,2,3,4\}$, we obtain the complete side information gain profile in Table IV.
TABLE IV: Minimum distance, message rate, and side information gain in each subcode $\mathscr{C}_{S}$.
Index set Minimum distance Message rate $\Gamma(\mathscr{C},S)$
$S=\{1\}$ $3$ $31/2\log_{2}$ 3 12.0412
$S=\{2\}$ $3$ $31/4\log_{2}$ 3 24.0824
$S=\{3\}$ $\sqrt{10}$ $51/4\log_{2}$ 5 17.2271
$S=\{4\}$ $\sqrt{10}$ $51/4\log_{2}$ 5 17.2271
$S=\{1,2\}$ $3$ $271/4\log_{2}$ 27 8.0274
$S=\{1,3\}$ $\sqrt{30}$ 1/$451/4\log_{2}$ 45 10.7586
$S=\{1,4\}$ $\sqrt{30}$ 1/$451/4\log_{2}$ 45 10.7586
$S=\{2,3\}$ $\sqrt{15}$ $151/4\log_{2}$ 15 12.0412
$S=\{2,4\}$ $\sqrt{15}$ $151/4\log_{2}$ 15 12.0412
$S=\{3,4\}$ $5$ $251/4\log_{2}$ 25 12.0412
$S=\{1,2,3\}$ $3\sqrt{5}$ $1351/4\log_{2}$ 135 9.3443
$S=\{1,2,4\}$ $3\sqrt{5}$ $1351/4\log_{2}$ 135 9.3443
$S=\{1,3,4\}$ $5\sqrt{3}$ 1/$451/4\log_{2}$ 45 9.5987
$S=\{2,3,4\}$ $5\sqrt{3}$ $751/4\log_{2}$ 75 12.0412
From Table IV we observe that the side information gain $\Gamma$ depends significantly on the specific subset of messages that are known. For example, knowing message $a_{2}$, i.e., $S=\{2\}$, yields $\Gamma\approx 24.08$ dB/bit/dim, while knowing $a_{1}$, i.e., $S=\{1\}$, results in $\Gamma\approx 12.04$ dB/bit/dim. Interestingly, when both messages are known, the side information gain $\Gamma$ reduces to $\approx 8.03$ dB/bit/dim. This behaviour results from the multilevel structure of Construction $\pi_{A}$, where each component code influences the minimum distance $d_{S}$, and suggest that the generator of each code must be chosen carefully. Such insights are valuable for future designs, as they suggest that prioritizing specific messages as side information can lead to more efficient decoding strategies. These properties illustrate that Construction $\pi_{A}$ is particularly well suited for networks where receivers have unequal access to side information.
Previous works have also explored the use of CRT for constructing lattice index codes [21, 42, 4]. However, those approaches typically restrict the code $\mathscr{C}$ to have size $q^{n}$, where $q$ is the product of the primes involved and $n$ is the real dimension of the space. Construction $\pi_{A}$ offers greater flexibility in index code design, allowing for code sizes of the form $|\mathscr{C}|=|\mathscr{C}_{1}^{(1)}|\cdot|\mathscr{C}_{1}^{(2)}|\cdot\ldots\cdot|\mathscr{C}_{r}^{(1)}|\cdot|\mathscr{C}_{r}^{(2)}|=p_{1}^{k_{1}^{(1)}+k_{1}^{(2)}}\cdot\ldots\cdot p_{r}^{k_{r}^{(1)}+k_{r}^{(2)}}$, where $1\leq k_{j}^{(i)}\leq n$ is the rank of the linear code $\mathscr{C}_{j}^{(i)}$, $i=1,2$ and $j=1,\ldots,r$. In the case of index coding using Construction $\pi_{A}$ lattices over the ring of integers a detailed investigation can be found in [57, 58].
VII Conclusion and Perspectives